OpenStudy (anonymous):

Please help! Caley has a bag that contains 9 blue marbles and 13 yellow marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Caley selects a blue marble and then a yellow marble? Round your answer to the nearest percent. P(blue and yellow) ≈ ______%

jimthompson5910 (jim_thompson5910):

what is the probability of picking a blue marble?

OpenStudy (anonymous):

Idk lol

jimthompson5910 (jim_thompson5910):

how many blue marbles are there? how many marbles are there total?

OpenStudy (anonymous):

What is the probability that Caley selects a blue marble and then a yellow marble?

jimthompson5910 (jim_thompson5910):

we'll get there, first answer these two subquestions how many blue marbles are there? how many marbles are there total?

OpenStudy (anonymous):

Oh sorry 9 and 22

jimthompson5910 (jim_thompson5910):

so the probability of picking blue is the fraction 9/22

jimthompson5910 (jim_thompson5910):

after one marble is taken out, there are 22-1 = 21 marbles left there are 13 yellow marbles which is why the probability of picking yellow, after you pick blue, is 13/21

jimthompson5910 (jim_thompson5910):

multiply these two fractions out (9/22)*(13/21) = ??

OpenStudy (anonymous):

39/54

OpenStudy (anonymous):

i ment 139/54

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(9/22)*(13/21) = 39/154 which is approximately 0.2532 convert 0.2532 to a percentage to get 25.32% then round this to the nearest percent to get 25%

OpenStudy (anonymous):

okay thank you!

jimthompson5910 (jim_thompson5910):

np